OpenStudy (anonymous):

Finding radius of convergence:

OpenStudy (anonymous):

\[\sum_{n=0}^{\infty} nx ^{n}\]

OpenStudy (anonymous):

I've tried the ratio test, and I ended up with divergence, which is wrong. I'm not sure where I messed up.

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

\[\lim_{n \rightarrow \infty} n\]

OpenStudy (anonymous):

In short.

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

I think I messed up somewhere after I got this: \[\lim_{n \rightarrow \infty} \frac{ (n+1)x }{ n }\]

OpenStudy (anonymous):

ok that looks good

OpenStudy (anonymous):

even better is \[\frac{n+1}{n}|x|\]

OpenStudy (zarkon):

need absolute value on x....;)

OpenStudy (anonymous):

Yeah, I have the absolute value, but I didn't wanna chance it with the equation editor. I'm not very good with it. lol.

OpenStudy (anonymous):

is it more or less obvious what \[\lim_{n\to \infty}\frac{n+1}{n}\] is?

OpenStudy (anonymous):

It's gonna be 1, correct?

OpenStudy (anonymous):

I feel like this is a lot easier than I made it.

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

yes, it is

OpenStudy (anonymous):

i mean yes it is one yes it is easy

OpenStudy (anonymous):

Ok, so we have that \[\lim_{n=0 \rightarrow \infty} |x|\]

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

hmmm

OpenStudy (anonymous):

Uh oh. I don't like "hmmm".

OpenStudy (anonymous):

oh i see what you are saying \[\lim_{n\to \infty}\frac{n+1}{n}|x|=|x|\] at least i hope that is what you meant

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

then that is right, and the answer now should be pretty clear it converges when the limit is less than one, so in this case \[|x|<1\]

OpenStudy (anonymous):

Oh. That's all I need to do? I thought I needed to test the intervals by plugging x=1 and -1 into the original equation in place of x?

OpenStudy (anonymous):

the radius of convergence is 1, that is \(|x|<1\) as for the interval of convergence you do need to test the endpoints to see if the interval is \[(-1,1)\\ [-1,1]\\ [-1,1)\\ (-1,1]\]

OpenStudy (anonymous):

When I tested the intervals, I got really messed up.

OpenStudy (anonymous):

that plugging is business is checking the endpoints not the interval

OpenStudy (anonymous):

i sense you are confused

OpenStudy (anonymous):

That's a wild understatement. lol.

OpenStudy (anonymous):

the interval you know it converges on is \((-1,1)\) i.e. \(|x|<1\)

OpenStudy (anonymous):

in english, (well math english) you would say "the radius of convergence is one"

OpenStudy (anonymous):

Ok. I gotcha. Don't let the name confuse you, I'm 100% english.

OpenStudy (anonymous):

now that you have established that fact, the only question that remains is does it converge at the endpoints (one or the other or both) or that interval

OpenStudy (anonymous):

lol by "in english " i meant "this is what you say in words, as opposed to symbols" not "this is how we say it here in the US"

OpenStudy (anonymous):

Oh, ok, lol. If I were to test the endpoints, would I just plug them into the original equation? Or how would I go about that?

OpenStudy (anonymous):

in any case i hope it is clear that your final job is to replace \(x\) by \(-1\) and by \(1\) to test the endpoints of the interval

OpenStudy (anonymous):

where you see an \(x\) put a \(1\)

OpenStudy (anonymous):

Yeah, ok, so it was much easier than I was making it. Thank you very much!

OpenStudy (anonymous):

yw that is where you got \[\sum n\] which surely does not converge that just means the right hand endpoint is not in the interval of convergence

OpenStudy (anonymous):

btw not that you asked, but this one has a nice an easy to find closed form

OpenStudy (anonymous):

Yeah, I have an exam tomorrow and I'm refreshing on this stuff. Based on this, I'm not so sure of how it's gonna go. I'm not expecting easy and nice closed forms.

OpenStudy (anonymous):

you know \[\sum x^n=\frac{1}{1-x}\] right?

OpenStudy (anonymous):

I didn't know that at all.

OpenStudy (anonymous):

oops

OpenStudy (anonymous):

Well, that may help tomorrow!

OpenStudy (anonymous):

for \(|x|<1\) it is true and a very useful fact to know

OpenStudy (anonymous):

i will bet you knew it if you every summed a geometric sequence you used it

OpenStudy (anonymous):

Oh wait, that's the sum for an infinite geometric series. I'm just used to seeing the summation with a starting value and ending in infinity.

OpenStudy (anonymous):

So yeah, I've used it. Just didn't recognize it right away.

OpenStudy (anonymous):

ok then take the derivative, get \[\sum nx^{n-1}=\frac{1}{(1-x)^2}\]

OpenStudy (anonymous):

multiply by \(x\) and get \[\sum nx^n=\frac{x}{(1-x)^2}\]

OpenStudy (anonymous):

Wow. That would have taken a bit less time. Might have to remember that for tomorrow.

OpenStudy (anonymous):

good luck!~